If the volume of an ideal gas is doubled while its temperature is quadrupled, does the pressure
(a) remain the same
(b) decreases by a factor of 2
(c) decrease by a factor of 4
(d) Increase by a factor of 2
(e) increase by a factor of 4
1
Expert's answer
2021-02-11T17:11:19-0500

"\frac{p_1V_1}{T_1}=\frac{p_2\cdot 2V_1}{4T_2},"

"p_2=2p_1,"

(d) increase by a factor of 2.